10. Determine the pH of a solution that is 0.20 M in NH3 and 0.3 M NH4Cl.
Answer #1

Kb = 1.8*10^-5

pKb = - log (Kb)

= - log(1.8*10^-5)

= 4.745

use:

pOH = pKb + log {[conjugate acid]/[base]}

= 4.745+ log {0.3/0.2}

= 4.921

use:

PH = 14 - pOH

= 14 - 4.9208

= 9.0792

Answer: 9.08